OpenStudy (anonymous):

The sum of two numbers is x. If one of the numbers is 12, what is the other? 12 - x x - 12 12 + x 12 x

OpenStudy (campbell_st):

well test the answer choices to see which gives than answer of x (a) 12 + 12 - x = 24 - x not x so (a) isn't the answer repeat the process

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

thanks

OpenStudy (campbell_st):

it's not C 12 + 12 + x = 24 + x does this equal the required number x..?

OpenStudy (anonymous):

good point i didnt do + x

OpenStudy (anonymous):

D is the anser